A body moves from a position r1 =i +2j -3k m and r2 =2i +3j -3k m under the influence of a constant force F 3i +I- k calculate workdone by force


1
Expert's answer
2021-07-19T09:50:29-0400
"\bold{r_2-r_1}=(2-1,3-2,-3+3)\\=(1,1,0)\ m\\
W=(3,1,-1)(1,1,0)=4\ J"